OpenStudy (anonymous):

using double angle identities, find the exact solutions for 0

OpenStudy (vishweshshrimali5):

Are you sure it is 5 cos2 in RHS ? @joshnel

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

NO sorry, it is 5cosx RHS

OpenStudy (vishweshshrimali5):

Yeah thought so :)

OpenStudy (vishweshshrimali5):

\[\large{3+\cos 2x = 5\cos x}\] First of all we will use this formula: \[\large{\cos 2x = 2\cos^2 x - 1}\]

OpenStudy (vishweshshrimali5):

We will get: \[\large{3+2\cos^2 x - 1 = 5\cos x}\] \[\large{\implies 2\cos^2 x + 2 - 5\cos x = 0}\]

OpenStudy (vishweshshrimali5):

Any doubt ?

OpenStudy (anonymous):

No, and now we factor?

OpenStudy (vishweshshrimali5):

Yeah

OpenStudy (vishweshshrimali5):

\[\large{2\cos^2 x - 4\cos x - \cos x + 2 = 0}\] \[\large{\implies 2\cos x (\cos x - 2) - (\cos x -2 ) = 0}\] \[\large{\implies (2\cos x - 1)(\cos x - 2) = 0}\]

OpenStudy (anonymous):

then set each factor to zero, then substituts cos=y, solve for y, then sub y=cos back into the equation?

OpenStudy (vishweshshrimali5):

Perfect !!!!!!!!! Just remember that cos x cannot be less than -1 and more than 1

OpenStudy (anonymous):

Suppose that it is? or is that just impossible?

OpenStudy (vishweshshrimali5):

\[\large{-1 \le \cos x \le 1}\] Thus, one solution would automatically be discarded as cos x = 2 is not a possible value

OpenStudy (anonymous):

wait, so the factor (cosx-2) is automatically disregarded?

OpenStudy (vishweshshrimali5):

Yup... you will have to consider only 2cos x - 1 = 0

OpenStudy (vishweshshrimali5):

Because cos x - 2= 0 will give cos x = 2 which is not possible

OpenStudy (anonymous):

Oh, i see, i literally just repeated what you typed, sorry about that.

OpenStudy (vishweshshrimali5):

Its okay

OpenStudy (anonymous):

so the only possible solution is cos = 1/2

OpenStudy (vishweshshrimali5):

Yup

OpenStudy (anonymous):

ok, thank you so much, i really appreciate the help. I think i have it down, its just applying the double angle formulas

OpenStudy (vishweshshrimali5):

Your welcome :)